A grandfather clock is designed so that one swing of the pendulum in either direction takes 1.00 s. What is the length of the pendulum?
1 : 0.993 m
2 : 0.248 m
3 : 0.062 m
4 : 0.100 m
5 : None of the above
1
Expert's answer
2020-08-02T15:07:46-0400

One period in a pendulum is 2 swings, therefore the period "T=2.00 s."

"T=\frac{1}{f}=2\pi\sqrt\frac{L}{g}"

After solving for "L" , we get e get;


"L=g(\frac{T}{2\pi})^2"


"=(9.80m/s^2)(\frac{2.00s}{2\pi})^2"


"=0.993 m"

Therefore, the correct answer is option "1 : 0.993 m"